# LLM Query Scheduling with Prefix Reuse and Latency Constraints

### v1

@@ 1 Introduction
The rapid integration of large language models (LLMs) into online systems has spurred significant research aimed at improving their inference efficiency. Unlike traditional batch-processing environments, online usage demands a nuanced understanding of performance, prioritizing what is often termed “goodput” [24]—the maximum number of requests that can be served while meeting stringent constraints on time-to-first-token (TTFT) and time-per-output-token (TPOT). This shift in focus underscores the necessity of developing advanced inference and serving algorithms that optimize goodput, ensuring efficient and cost-effective deployment of LLMs in latency-sensitive applications.
An important approach to improving LLM inference efficiency is prefix reuse. In autoregressive LLMs, prompts with shared prefixes can leverage the intermediate representations (i.e. key and value tensors, often stored in the key-value (KV) cache) of the common prefix, reducing both memory usage and computational overhead during inference. A notable implementation of this idea is RadixAttention, which stores the KV cache of processed queries and automatically reuses shared prefixes by maintaining a radix tree [23]. This tree tracks cached values and matches them with new incoming prompts, enabling efficient reuse without requiring manual intervention. To manage memory constraints imposed by the fixed size of the KV cache, the system employs a least recently used (LRU) eviction policy for cached values. This method is particularly impactful in scenarios where many prompts share prefixes, offering substantial savings in compute and memory. Moreover, its “out-of-the-box” functionality simplifies adoption by eliminating the need for users to manually analyze or encode prefix structures, making it a practical solution for real-world applications.
@@
The main scheduling algorithms considered in prior work are First Come First Serve (FCFS), which processes queries in arrival order, and Longest Prefix Match (LPM) [^1] LPM is the default query scheduling algorithm in SGLang v0.4.1. Meanwhile, FCFS is an option in SGLang and the default scheduling algorithm in vLLM v0.6.6., which greedily maximizes KV cache reuse at each scheduling step. In the offline setting (where all queries are known in advance), Theorem 3.1 in [23] shows that LPM indeed maximizes cache reuse. However, the online setting, where queries arrive over time under tight time-to-first-token constraints, was not characterized.
In this work, we provide a theoretical exploration of RadixAttention scheduling in the online regime, focusing on non-preemptive scheduling. Our analysis and experiments show that existing scheduling approaches can lead to large TTFT spikes under heavy traffic, motivating a new scheduling algorithm that more robustly balances prefix reuse and waiting time. Specifically, we design a mechanism that exploits the benefits of LPM but mitigates its performance risks in high-load scenarios, resulting in superior performance for long-prompt, short-output (i.e., prefill-dominant) queries where RadixAttention delivers significant efficiency gains. Examples of such applications include document summarization, coding assistants, and prompts with detailed instruction sets among others [2]. Recently, a well-studied generation use case that can benefit from faster prefill is the scaling of test-time compute (TTC) via best-of-N sampling, where $N$ outputs are generated in parallel and a verifier model is used to score them [15, 3]. TTC has enabled recent LLMs to dramatically improve performance on reasoning benchmarks for domains like coding and math [7].
@@ 1.1 Contributions
  • In Section 2, we introduce a formal model for analyzing the LLM query scheduling problem. Our model draws inspiration from prior work on the roofline model [8] and incorporates experimental observations to ensure practical relevance. Despite its grounding in real-world considerations, the model remains sufficiently simple to facilitate analytical insights, including a formal specification of the “query stream” (Definition 1) and a computational model for LLMs (Definition 2), enabling further exploration of the problem.
  • We show that the decision problem of determining whether a TTFT constraint can be satisfied for a given query stream is NP-Hard when using Radix Attention (Theorem 1). This is in contrast to the case without Radix Attention, where latency is trivially minimized by FCFS, or the case with Radix Attention and uniform arrival times, where latency is minimized by LPM [23].
@@
  • Finally, in Appendix A we prove that an approximation algorithm exists that, for a length $n$ query stream $\mathcal{Q}$ and TTFT constraint $T$, either 1) certifies no schedule exists satisfying the TTFT constraint, or, 2) returns a schedule such that the $(1-p)$-th percentile TTFT is at most $T$ for $p\in(0,1)$. Furthermore, this algorithm runs in $\mathcal{O}(n\cdot\exp(1/p\log 1/p))$ time.
@@ 1.2 Related Work
There has been significant interest in prior work on scheduling LLM queries in order to maximize throughput while satisfying TTFT constraints. Some examples are FastSwitch, which employs a priority-based scheduler with preemption to dynamically allocate resources to effectively reduce TTFT and GPU idleness [14], and Orca, which employs iteration-level scheduling, processing each model iteration separately allowing for flexible batching and immediate response to newly arrived requests [20]. Many other papers increase throughput subject to latency constraints by innovations to the processing schedule of incoming prompts [24, 11, 9, 1]. Collectively, this prior work underscores the importance of dynamic scheduling in achieving high throughput without compromising latency guarantees.
The most relevant work to ours addresses scheduling of LLM queries with consideration of prefix reuse. Zheng et al. [23] introduced RadixAttention and the LPM scheduler, which we build upon, but their work provides only limited exploration of scheduling strategies. Srivatsa et al. [16] proposes a priority-based local scheduler in Section 3.3 aimed at balancing prefix reuse and waiting times. The scheduler works by assigning requests to priority groups based on prefix cache hit rate and then selects a number of prompts in each group proportionally to its priority, but no accompanying analysis is provided. Qin et al. [12] integrates RadixAttention in local instances but primarily emphasizes maintaining a disaggregated KV cache to balance the decode and prefill phases, enhancing throughput while adhering to latency constraints in highly overloaded scenarios.
Analytic exploration of LLM inference efficiency is relatively limited, with some notable examples. Kim et al. [10] propose INFERMAX, an analytical framework for evaluating schedulers and deriving theoretical performance bounds, while identifying prefix-sharing awareness as a future direction. Yang et al. [19] analyze the behavior of LLM queries using an M/G/1 queue while accounting for unknown decoding length, and Guldogan et al. [6] examine multi-bin batching to boost throughput. These studies underscore the value of theoretical approaches, which our work advances through a focus on prefix-aware scheduling.
@@ 2 Computational Model of RadixAttention
In this section, we provide a model of LLM computation that allows for theoretical study of RadixAttention under different scheduling algorithms. Consider a single LLM instance (that may be on a single GPU or parallelized across multiple GPUs in various ways). This LLM instance can process queries using continuous batching. The time for a single pass can be understood from the “roofline model” (see Appendix A.2 of [8] for details on applying the roofline model to LLM inference).
@@
Consider a stream of queries $\mathcal{Q}=\{(\mathbf{x}_{i},t_{i})\}_{i\in[n]}$. Suppose these queries are processed in the order $\mathbf{x}_{j_{1}},\mathbf{x}_{j_{2}},\ldots,\mathbf{x}_{j_{n}}$. Let $R(j_{k})$ denote the completion time of the $j_{k}$-th query. Then
$R$ $(j_{1})=\lvert\mathbf{x}_{j_{1}}\rvert,$
@@ 2.1 Batch Size and Prefix Cache
The original specification of RadixAttention does not leverage reuse of prefixes of prompts in a single batch (see Algorithm 1 in [23]). However, more recent development in SGLang has enabled within batch prefix sharing. In this case, behavior of scheduling algorithms in the batched setting can be closely approximated by the batch size one setting in Definition 2. Computing a query processing order and then dividing the queries into $B$ sized bins sequentially provides a schedule for the $B$ batch size setting that attains the same amount of prefix reuse. The only difference is that the time that the $i$-th query is finished being processed will be the same as when the $\lceil i/B\rceil$-th query is finished being processed. This difference will be negligible when the query arrival rate is high relative to the batch size.
@@ 3.1 NP-Hardness of Feasibility Determination
We provide the definition of the 3-PARTITION problem, which is NP-Hard in the strong sense. See SP15 in the appendix of [5] for reference.
@@
Set the overall query stream to be
$\mathcal{Q}\;=\;\mathcal{X}\,\cup\,\mathcal{Y}\,\cup\,\mathcal{Z}\,\cup\,\{\mathbf{w}_{1},\mathbf{w}_{2}\}.$
Scheduling constraints.
Observe:
@@
  • The total time to process $\mathcal{Y}$ and $\mathcal{Z}$ (accounting for maximum $\mathbf{y}_{i}$$\mathbf{z}_{i}$ overlap) is:
    $\sum_{\mathbf{y}_{i}\in\mathcal{Y}}|\mathbf{y}_{i}|\;=\;m^{2}H\;=\;T-mH.$
@@
Then $\mathbf{y}_{i+1}$ must be processed starting from $T+i(H+mH)+H$ in order to finish by $T+(i+1)(H+mH)$ and maintain the “no idle time” schedule. Since $\mathbf{z}_{i}$ is identical to $\mathbf{y}_{i}$, its processing time is zero in this case. Therefore, between $\mathbf{z}_{i}$ finishing at $\,T+i(H+mH)$ and $\mathbf{y}_{i+1}$ starting at $\,T+i(H+mH)+H,$ there is exactly a length-$H$ sub-interval available. Since $\sum_{\mathbf{x}_{i}\in\mathcal{X}}|\mathbf{x}_{i}|=mH,$ the only way to fill these $m$ sub-intervals continuously is to divide $\{\mathbf{x}_{i}\}$ into $m$ disjoint groups, each summing to exactly $H$. But deciding such a partition is precisely the 3-PARTITION problem. Consequently, scheduling $\mathcal{Q}$ to meet the TTFT constraint is possible if and only if the instance $\mathcal{A}$ of 3-PARTITION admits a feasible partition.
In Appendix A, we further explore the general problem of scheduling queries to satisfy a TTFT constraint. In Theorem 4, we prove that there exists an algorithm that accepts a query stream and TTFT constraint $T$ and either certifies that there is no schedule satisfying constraint $T$ or returns a schedule such that the $(1-p)$-th percentile TTFT is at most $T$ in $\mathcal{O}(n\cdot\exp(1/p\log 1/p))$ time.
@@ 4.1 Data Generative Model
In this section, we restrict our attention to instances where queries in a queue approximately follow a tree structure of low height. An example of such prompt structure used in [17] is:
[…]
Here, all queries share the same (base_prompt), and multiple queries may share the same (user_context).
Examples of such structures include generative usecases: personalized content generation [22], conversational context-aware question answering [21], and the predictive usecase with LLMs as engagement predictors in recommendation systems [18]. These scenarios exemplify applications where the prompt structure remains consistent while the user context varies, allowing for efficient processing and relevant responses. By focusing on such structured instances, we can better analyze and optimize the computational models for LLMs under constrained scheduling conditions.
@@ 4.2 TTFT Improvement from $k$-LPM
In scenarios where each query can be processed swiftly—specifically, when the processing time of a given query is less than the inter-arrival interval $s$—the FCFS scheduling algorithm is optimal. However, in more practically relevant burst-traffic regimes where queries arrive in rapid succession and cannot be processed fast enough, a backlog of unprocessed queries inevitably forms. To illustrate this, consider a toy example with $c_{\mathrm{attn}}=0$, $n=4$ queries, a replication factor $k=2$, and parameters $u=5$ and $d=5$. The queries are denoted as $\mathbf{x}_{1}=\texttt{(user)}_{1}\texttt{(doc)}_{1}$, $\mathbf{x}_{2}=\texttt{(user)}_{2}\texttt{(doc)}_{2}$, $\mathbf{x}_{3}=\texttt{(user)}_{1}\texttt{(doc)}_{3}$, and $\mathbf{x}_{4}=\texttt{(user)}_{2}\texttt{(doc)}_{4}$. In the case where the inter-arrival time $s=10$, FCFS scheduling is clearly optimal, resulting in a uniform TTFT of 10 units for each query. Conversely, in the case with $s=0$, representing a burst-traffic scenario, the LPM scheduling strategy becomes optimal. Under FCFS, the processing order is $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$, $\mathbf{x}_{4}$ and $\texttt{TTFT}_{i}=10\cdot i$ with $\max(\texttt{TTFT}_{i})=\texttt{TTFT}_{4}=40$. Under LPM, the processing order is rearranged to $\mathbf{x}_{1}$, $\mathbf{x}_{3}$, $\mathbf{x}_{2}$, $\mathbf{x}_{4}$ with TTFTs of 10, 15, 25, and 30, and $\max(\texttt{TTFT}_{i})=30$. This example underscores the improvement of TTFT from cache reuse in practical scenarios.
@@
Intuitively, the theorem shows that when $s$ is small and $u$ is large, LPM is much better than FCFS, and $k$-LPM retains this advantage. On the other hand, when $s$ is relatively large, FCFS is better and $k$-LPM retains a $\frac{1}{k}$ factor of the $sn$ reduction in TTFT. For intermediate values of $s$ and $u$, $k$-LPM is better than both algorithms as we prove in Corollary 3.
@@
  • (LPM) For every $\epsilon>0$ and $\delta\in(0,1)$, there is an $n_{0}$ such that for all $n\geq n_{0}$, with probability at least $1-\delta$ (with respect to random shuffle and randomness in LPM),
    $\max_{i\in[n]}\,\texttt{TTFT}_{i}~{}\;\geq\;T\;+\;\bigl{(}1-\epsilon\bigr{)}\;n\,\Bigl{(}\tfrac{u}{k}+d\Bigr{)}.$
  • (FCFS) For every $\epsilon>0$ and $\delta\in(0,1)$, there is an $n_{0}$ such that for all $n\geq n_{0}$, with probability at least $1-\delta$ (over the random shuffle),
@@
  • ($k$-LPM) Deterministically (i.e. for any arrival order), Algorithm 1 satisfies
    $\max_{i\in[n]}\,\texttt{TTFT}_{i}~{}\;\leq\;T\;+\;n\,\Bigl{(}\tfrac{u}{k}+d-\tfrac{s}{k}\Bigr{)}.$

### v3

@@ 1 Introduction
The rapid integration of large language models (LLMs) into online systems has spurred significant research aimed at improving their inference efficiency. Unlike traditional batch-processing environments, online usage demands a nuanced understanding of performance, prioritizing what is often termed “goodput” zhong2024distserve—the maximum number of requests that can be served while meeting stringent constraints on time-to-first-token (TTFT) and time-per-output-token (TPOT). This shift in focus underscores the necessity of developing advanced inference and serving algorithms that optimize goodput, ensuring efficient and cost-effective deployment of LLMs in latency-sensitive applications.
An important approach to improving LLM inference efficiency is prefix reuse. In autoregressive LLMs, prompts with shared prefixes can leverage the intermediate representations (i.e. key and value tensors, often stored in the key-value (KV) cache) of the common prefix, reducing both memory usage and computational overhead during inference. A notable implementation of this idea is RadixAttention, which stores the KV cache of processed queries and automatically reuses shared prefixes by maintaining a radix tree zheng2024sglang. This tree tracks cached values and matches them with new incoming prompts, enabling efficient reuse without requiring manual intervention. To manage memory constraints imposed by the fixed size of the KV cache, the system employs a least recently used (LRU) eviction policy for cached values. This method is particularly impactful in scenarios where many prompts share prefixes, offering substantial savings in compute and memory. Moreover, its “out-of-the-box” functionality simplifies adoption by eliminating the need for users to manually analyze or encode prefix structures, making it a practical solution for real-world applications.
@@
The main scheduling algorithms considered in prior work are First Come First Serve (FCFS), which processes queries in arrival order, and Longest Prefix Match (LPM) [^1] LPM is the default query scheduling algorithm in SGLang v0.4.1. Meanwhile, FCFS is an option in SGLang and the default scheduling algorithm in vLLM v0.6.6., which greedily maximizes KV cache reuse at each scheduling step. In the offline setting (where all queries are known in advance), Theorem 3.1 in zheng2024sglang shows that LPM indeed maximizes cache reuse. However, the online setting, where queries arrive over time under tight time-to-first-token constraints, was not characterized.
In this work, we provide a theoretical exploration of RadixAttention scheduling in the online regime, focusing on non-preemptive scheduling. Our analysis and experiments show that existing scheduling approaches can lead to large TTFT spikes under heavy traffic, motivating a new scheduling algorithm that more robustly balances prefix reuse and waiting time. Specifically, we design a mechanism that exploits the benefits of LPM but mitigates its performance risks in high-load scenarios, resulting in superior performance for long-prompt, short-output (i.e., prefill-dominant) queries where RadixAttention delivers significant efficiency gains. Examples of such applications include document summarization, coding assistants, and prompts with detailed instruction sets among others anthropic_prompt_caching. Recently, a well-studied generation use case that can benefit from faster prefill is the scaling of test-time compute (TTC) via best-of-N sampling, where $N$ outputs are generated in parallel and a verifier model is used to score them snell2024scaling; cobbe2021training. TTC has enabled recent LLMs to dramatically improve performance on reasoning benchmarks for domains like coding and math guo2025deepseek.
@@ 1.1 Contributions
  • In Section 2, we introduce a formal model for analyzing the LLM query scheduling problem. Our model draws inspiration from prior work on the roofline model imai2024predicting and incorporates experimental observations to ensure practical relevance. Despite its grounding in real-world considerations, the model remains sufficiently simple to facilitate analytical insights, including a formal specification of the “query stream” (Definition 1) and a computational model for LLMs (Definition 2), enabling further exploration of the problem.
  • We show that the decision problem of determining whether a TTFT constraint can be satisfied for a given query stream is NP-Hard when using Radix Attention (Theorem 1). This is in contrast to the case without Radix Attention, where latency is trivially minimized by FCFS, or the case with Radix Attention and uniform arrival times, where latency is minimized by LPM zheng2024sglang.
@@
  • Finally, in Appendix A we prove that an approximation algorithm exists that, for a length $n$ query stream $\mathcal{Q}$ and TTFT constraint $T$, either 1) certifies no schedule exists satisfying the TTFT constraint, or, 2) returns a schedule such that the (1p)(1-p)-th percentile TTFT is at most $T$ for $p\in(0,1)$. Furthermore, this algorithm runs in $\mathcal{O}(n\cdot\exp(1/p\log 1/p))$ time.
@@ 1.2 Related work
There has been significant interest in prior work on scheduling LLM queries in order to maximize throughput while satisfying TTFT constraints. Some examples are FastSwitch, which employs a priority-based scheduler with preemption to dynamically allocate resources to effectively reduce TTFT and GPU idleness shen2024fastswitch, and Orca, which employs iteration-level scheduling, processing each model iteration separately allowing for flexible batching and immediate response to newly arrived requests yu2022orca. Many other papers increase throughput subject to latency constraints by innovations to the processing schedule of incoming prompts zhong2024distserve; patel2024splitwise; jain2024intelligent; agrawal2024taming. Collectively, this prior work underscores the importance of dynamic scheduling in achieving high throughput without compromising latency guarantees.
The most relevant work to ours addresses scheduling of LLM queries with consideration of prefix reuse. zheng2024sglang introduced RadixAttention and the LPM scheduler, which we build upon, but their work provides only limited exploration of scheduling strategies. srivatsa2024preble proposes a priority-based local scheduler in Section 3.3 aimed at balancing prefix reuse and waiting times. The scheduler works by assigning requests to priority groups based on prefix cache hit rate and then selects a number of prompts in each group proportionally to its priority, but no accompanying analysis is provided. qin2024mooncake integrates RadixAttention in local instances but primarily emphasizes maintaining a disaggregated KV cache to balance the decode and prefill phases, enhancing throughput while adhering to latency constraints in highly overloaded scenarios.
Analytic exploration of LLM inference efficiency is relatively limited, with some notable examples. kim2024effect propose INFERMAX, an analytical framework for evaluating schedulers and deriving theoretical performance bounds, while identifying prefix-sharing awareness as a future direction. yang2024queueing analyze the behavior of LLM queries using an M/G/1 queue while accounting for unknown decoding length, and guldogan2024multi examine multi-bin batching to boost throughput. These studies underscore the value of theoretical approaches, which our work advances through a focus on prefix-aware scheduling.
@@ 2 Computational model of RadixAttention
In this section, we provide a model of LLM computation that allows for theoretical study of RadixAttention under different scheduling algorithms. Consider a single LLM instance (that may be on a single GPU or parallelized across multiple GPUs in various ways). This LLM instance can process queries using continuous batching. The time for a single pass can be understood from the “roofline model” (see Appendix A.2 of imai2024predicting for details on applying the roofline model to LLM inference).
@@
Consider a stream of queries $\mathcal{Q}=\{(\mathbf{x}_{i},t_{i})\}_{i\in[n]}$. Suppose these queries are processed in the order $\mathbf{x}_{j_{1}},\mathbf{x}_{j_{2}},\ldots,\mathbf{x}_{j_{n}}$. Let $R(j_{k})$ denote the completion time of the $j_{k}$-th query. Then $R(j_{1})=\lvert\mathbf{x}_{j_{1}}\rvert+c_{\mathrm{attn}}\lvert\mathbf{x}_{j_{1}}\rvert^{2}$, and,
@@ 2.1 Batch size and prefix cache
The original specification of RadixAttention does not leverage reuse of prefixes of prompts in a single batch (see Algorithm 1 in zheng2024sglang). However, more recent development in SGLang has enabled within batch prefix sharing. In this case, behavior of scheduling algorithms in the batched setting can be closely approximated by the batch size one setting in Definition 2. Computing a query processing order and then dividing the queries into $B$ sized bins sequentially provides a schedule for the $B$ batch size setting that attains the same amount of prefix reuse. The only difference is that the time that the $i$-th query is finished being processed will be the same as when the $\lceil i/B\rceil$-th query is finished being processed. This difference will be negligible when the query arrival rate is high relative to the batch size.
@@ 3.1 NP-hardness of feasibility determination
We provide the definition of the 3-PARTITION problem, which is NP-Hard in the strong sense. See SP15 in the appendix of garey1979computers for reference.
@@
Set the overall query stream to be $\mathcal{Q}\;=\;\mathcal{X}\,\cup\,\mathcal{Y}\,\cup\,\mathcal{Z}\,\cup\,\{\mathbf{w}_{1},\mathbf{w}_{2}\}$. Observe:
@@
  • The total time to process $\mathcal{Y}$ and $\mathcal{Z}$ (accounting for maximum $\mathbf{y}_{i}$$\mathbf{z}_{i}$ overlap) is: $\sum_{\mathbf{y}_{i}\in\mathcal{Y}}|\mathbf{y}_{i}|\;=\;m^{2}H\;=\;T-mH.$
@@
Then $\mathbf{y}_{i+1}$ must be processed starting from $T+i(H+mH)+H$ in order to finish by $T+(i+1)(H+mH)$ and maintain the “no idle time” schedule. Since $\mathbf{z}_{i}$ is identical to $\mathbf{y}_{i}$, its processing time is zero in this case. Therefore, between $\mathbf{z}_{i}$ finishing at T+i(H+mH)\,T+i(H+mH) and $\mathbf{y}_{i+1}$ starting at T+i(H+mH)+H,\,T+i(H+mH)+H, there is exactly a length-$H$ sub-interval available. Since $\sum_{\mathbf{x}_{i}\in\mathcal{X}}|\mathbf{x}_{i}|=mH,$ the only way to fill these $m$ sub-intervals continuously is to divide $\{\mathbf{x}_{i}\}$ into $m$ disjoint groups, each summing to exactly $H$. But deciding such a partition is precisely the 3-PARTITION problem. Consequently, scheduling $\mathcal{Q}$ to meet the TTFT constraint is possible if and only if the instance $\mathcal{A}$ of 3-PARTITION admits a feasible partition. ∎
In Appendix A, we further explore the general problem of scheduling queries to satisfy a TTFT constraint. In Theorem 4, we prove that there exists an algorithm that accepts a query stream and TTFT constraint $T$ and either certifies that there is no schedule satisfying constraint $T$ or returns a schedule such that the (1p)(1-p)-th percentile TTFT is at most $T$ in $\mathcal{O}(n\cdot\exp(1/p\log 1/p))$ time.
@@ 4.1 Data generative model
In this section, we restrict our attention to instances where queries in a queue approximately follow a tree structure of low height. An example of such prompt structure used in 360brew is $\mathbf{x}=\texttt{(base\_prompt)(user\_context)(doc)}$. Here, all queries share the same (base_prompt), and multiple queries may share the same (user_context).
Examples of such structures include generative usecases: personalized content generation zhang2024personalization, conversational context-aware question answering zaib2022conversational, and the predictive usecase with LLMs as engagement predictors in recommendation systems wu2024survey. These scenarios exemplify applications where the prompt structure remains consistent while the user context varies, allowing for efficient processing and relevant responses. By focusing on such structured instances, we can better analyze and optimize the computational models for LLMs under constrained scheduling conditions.
@@ 4.2 TTFT improvement from $k$-LPM
In scenarios where each query can be processed swiftly—specifically, when the processing time of a given query is less than the inter-arrival interval ss—the FCFS scheduling algorithm is optimal. However, in more practically relevant burst-traffic regimes where queries arrive in rapid succession and cannot be processed fast enough, a backlog of unprocessed queries inevitably forms. To illustrate this, consider a toy example with $c_{\mathrm{attn}}=0$, $n=4$ queries, a replication factor $k=2$, and parameters $u=5$ and $d=5$. The queries are denoted as $\mathbf{x}_{1}=\texttt{(user)}_{1}\texttt{(doc)}_{1}$, $\mathbf{x}_{2}=\texttt{(user)}_{2}\texttt{(doc)}_{2}$, $\mathbf{x}_{3}=\texttt{(user)}_{1}\texttt{(doc)}_{3}$, and $\mathbf{x}_{4}=\texttt{(user)}_{2}\texttt{(doc)}_{4}$. In the case where the inter-arrival time $s=10$, FCFS scheduling is clearly optimal, resulting in a uniform TTFT of 10 units for each query. Conversely, in the case with $s=0$, representing a burst-traffic scenario, the LPM scheduling strategy becomes optimal. Under FCFS, the processing order is $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$, $\mathbf{x}_{4}$ and $\texttt{TTFT}_{i}=10\cdot i$ with $\max(\texttt{TTFT}_{i})=\texttt{TTFT}_{4}=40$. Under LPM, the processing order is rearranged to $\mathbf{x}_{1}$, $\mathbf{x}_{3}$, $\mathbf{x}_{2}$, $\mathbf{x}_{4}$ with TTFTs of 10, 15, 25, and 30, and $\max(\texttt{TTFT}_{i})=30$. This example underscores the improvement of TTFT from cache reuse in practical scenarios.
@@
Intuitively, the theorem shows that when $s$ is small and $u$ is large, LPM is much better than FCFS, and $k$-LPM retains this advantage. On the other hand, when $s$ is relatively large, FCFS is better and $k$-LPM retains a $\frac{1}{k}$ factor of the $sn$ reduction in TTFT. For intermediate values of $s$ and $u$, $k$-LPM is better than both algorithms as we prove in Corollary 3. Proofs for these results are provided in Appendix B.
@@
  • (LPM) For every $\epsilon>0$ and $\delta\in(0,1)$, there is an $n_{0}$ such that for all $n\geq n_{0}$, with probability at least $1-\delta$ (with respect to random shuffle and randomness in LPM): $\max_{i\in[n]}\,\texttt{TTFT}_{i}~\;\geq\;T\;+\;\bigl(1-\epsilon\bigr)\;n\,\Bigl(\tfrac{u}{k}+d\Bigr).$
  • (FCFS) For every $\epsilon>0$ and $\delta\in(0,1)$, there is an $n_{0}$ such that for all $n\geq n_{0}$, with probability at least $1-\delta$ (over the random shuffle):
@@
  • ($k$-LPM) Deterministically (i.e. for any arrival order), Algorithm 1 satisfies: $\max_{i\in[n]}\,\texttt{TTFT}_{i}~\;\leq\;T\;+\;n\,\Bigl(\tfrac{u}{k}+d-\tfrac{s}{k}\Bigr).$
Corollary 3.
For any values of $s$, $k$, $u$, and $d$ such that $0<s<u$ and $k>1$, and for any $\delta\in(0,1)$, there exists $n_{0}\in\mathbb{N}$ such that $k$-LPM achieves a lower maximum TTFT that LPM and FCFS simultaneously on the regular arrival shuffled queue (Definition 4) for any value of $n\geq n_{0}$ with probability at least $1-\delta$. This result holds for any value of $c_{\mathrm{attn}}\geq 0$ in the computational model of Definition 2.
5 Experiments
In our experiments, we use the Llama-3.1-8B-Instruct model dubey2024llama with tensor parallelism across eight A100 GPUs. We run the experiment using the SGLang v0.4.1 serving framework. In particular, we evaluate the timing metrics using SGLang’s serving benchmark utility sglang_bench_serving and only modify the benchmarking dataset. We implement the $k$-LPM algorithm as an extension to the current LPM implementation in SGLang. Finally, we construct the dataset used for benchmarking by sampling four prompts with shared user history from the 8k context length prompts described in 360brew for $2100$ prompts in total. We follow the textual interface of 360brew—(Instruction, Member Profile, Past Interactions, Question)—where the first three constitute the shared prefix and only the Question varies; see 360Brew, Table 1 for an example of this exact layout. We then randomly shuffle the ordering of these prompts and use $100$ for warm up of the benchmarking server.
5.1 Performance of $k$-LPM
We measure P99 TTFT vs. request rate for varying values of $k$ in the $k$-LPM algorithm. Our key observation (Figure 2) is that setting the hyperparameter to $k=2$ achieves reduced P99 TTFT across a wide range of request rates compared to FCFS and LPM. Note that SGLang’s benchmarking utility uses a Poisson arrival process, so “request rate” refers to the average number of requests per second. Additional experiment results are provided in Appendix C showing other TTFT percentile metrics and performance when varying the amount of prefix reuse.
Figure 2: We measure P99 TTFT versus request rate for five values of the hyperparameter $k$ on 2000 randomly shuffled prompts from the usecase described in 360brew. Note that $k=1$ corresponds to FCFS and $k=\infty$ corresponds to LPM.
  • $k$-LPM performance remains robust under realistic models and prompt data, even when theoretical assumptions are relaxed. Despite using a real LLM, an off-the-shelf serving framework, and an industrial prompt dataset, the observed scheduling behavior closely matches our theory. Here, prefix overlap structure is only approximately tree-like and random tie-breaking holds in practice; streaming processing (starting on first arrival) shows the $T\geq sn$” assumption is inconsequential; and even when $k$ differs from the true replica count (we used four replicas), choices like $k=2$ or $3$ still yield substantial P99 TTFT improvements.
6 Future work
The primary objectives of this work were to formalize the LLM query scheduling problem in the context of RadixAttention and to develop a practical scheduling algorithm for latency-sensitive applications. This theoretical framework not only informs the design of new methods, but is particularly valuable given the current lack of empirical benchmarks for comparing scheduling strategies. However, further advancements in this area will necessitate empirical evaluation using real-world arrival patterns. In particular, the described behavior depends on the complex interaction between constant factors dictating the arrival rate, query processing rate, and average prefix reuse, along with batch size, radix tree memory limit, and other factors in real serving frameworks. Greater understanding of LLM query scheduling in real systems necessitates a standardized, measurement-driven evaluation suite to characterize the practically relevant regime.
The proposed problem formulation and theoretical results leave many interesting extensions open. One open question is whether there is an algorithm which returns a schedule satisfying a constraint on the (1p)(1-p)-th percentile TTFT in $\operatorname{poly}(1/p)$ time, i.e., a polynomial time approximation scheme. Additional directions of interest would be extending the computational model of Section 2 to handle distributional query streams or non-constant decoding length. Finally, generalizing the data generative model of Definition 4 may be interesting to capture other properties of real data.
Appendix A Percentile TTFT constraint approximation
Theorem 4 proves there is an algorithm that runs in polynomial time with respect to the query stream length that either returns a schedule satisfying constraint $T$ on the (1p)(1-p)-th percentile TTFT or certifies that no feasible schedules exists such that all queries satisfy the TTFT constraint $T$. Note that there is a necessary gap where the the positive case pertain the the percentile TTFT problem and the negative case certifies no schedule satisfies a constraint on the maximum TTFT. This gap is fundamental, since otherwise we could solve the problem for the maximum TTFT in by adding enough unsatisfiable queries and solving the percentile TTFT decision problem.
Theorem 4.
  1. 1.
  2. 2.
    Returns a schedule for $\mathcal{Q}$ such that the (1p)(1-p)-th percentile TTFT is at most $T$,